OpenStudy (anonymous):

Plz help!!! :] Solve for t : 4t^3=108 Plzz and thankss!!!

OpenStudy (anonymous):

divide by 4, take the cubed root

OpenStudy (anonymous):

108/4 = 27 = t^3 cbrt{27} = 3 = t

OpenStudy (anonymous):

^sO THE answer is 3?

OpenStudy (anonymous):

Thank yuhh!!! ;]